# long-var-wrap = dv

### v2

@@ Appendix A Linear stability of $\bar{\varphi}=0$ solution
We test the stability of the solution presented in subsection IV.2 by inserting $\varphi_{a}(t,\bm{x})=\varphi_{a}(t,\bm{x};\bm{k})+\delta\varphi_{a}(\bm{x})$ in the equations of motion and expanding to first order in $\delta\varphi_{a}$. We note that, to linear order
$(\varphi+\delta\varphi)^{2}(\varphi_{a}+\delta\varphi_{a})=A^{2}\varphi_{a}+(A^{2}\delta_{ab}+2\varphi_{a}\varphi_{b})\delta\varphi_{b}.$ (23)
Using this, we write down the linear equation of motion
$\partial_{t}\delta\varphi_{a}=\Gamma\nabla^{2}\left[(r-\nabla^{2}+uA^{2})\delta_{ab}+2u\varphi_{a}\varphi_{b}+\alpha\epsilon_{ab}\right]\delta\varphi_{b}.$ (24)
In contrast to the homogenous solution, the overall operator $\nabla^{2}$ will now act on $\varphi$. We can expand the non-linear term using the chain rule. For an arbitrary function $f(x)$, if we define $A^{2}\mathcal{O}_{ab}f(x)\equiv\nabla^{2}[(A^{2}\delta_{ab}+2\varphi_{a}\varphi_{b})f(x)]$, we have
$A^{2}\mathcal{O}_{ab}=2\bigg{[}(\nabla^{2}\varphi_{a})\varphi_{b}+\varphi_{a}(\nabla^{2}\varphi_{b})+2(\bm{\nabla}\varphi_{a})\cdot(\bm{\nabla}\varphi_{b})$
$+2\Big{(}(\bm{\nabla}\varphi_{a})\varphi_{b}+\varphi_{a}(\bm{\nabla}\varphi_{b})\Big{)}\cdot\bm{\nabla}+\left(\frac{A^{2}\delta_{ab}}{2}+\varphi_{a}\varphi_{b}\right)\nabla^{2}\bigg{]}.$
As $\varphi_{1}\propto\sin\phi$ and $\varphi_{2}\propto\cos\phi$, we have the following properties
$\bm{\nabla}\varphi_{1}=\bm{k}\varphi_{2},\quad\bm{\nabla}\varphi_{2}=-\bm{k}\varphi_{1},\quad\bm{\nabla}^{2}\varphi_{a}=-k^{2}\varphi_{a}.$ (25)
We can therefore write $\bm{\nabla}\varphi_{a}=\bm{k}\epsilon_{ab}\varphi_{b}$, and so in Fourier space where $\bm{\nabla}\rightarrow-i\bm{q}$, the operator is
$A\mathcal{O}_{ab}$ $=-4k^{2}(\varphi_{a}\varphi_{b}+\epsilon_{ac}\epsilon_{bd}\varphi_{c}\varphi_{d})$ (26)
$\quad-4i\bm{k}\cdot\bm{q}\left(\varphi_{a}\epsilon_{bd}\varphi_{d}+\epsilon_{ac}\varphi_{c}\varphi_{b}\right)$
$\quad-q^{2}(A^{2}\delta_{ab}-2\varphi_{a}\varphi_{b}).$
We notice that $\mathcal{O}$ is a symmetric matrix. Writing out the components explicitly, we find
$\mathcal{O}_{11}$ $=4k^{2}-4i\bm{k}\cdot\bm{q}\sin 2\phi+q^{2}(1+2\cos^{2}\phi),$ (27)
$\mathcal{O}_{22}$ $=4k^{2}+4i\bm{k}\cdot\bm{q}\sin 2\phi+q^{2}(1+2\sin^{2}\phi),$ (28)
$\mathcal{O}_{12}$ $=4i\bm{k}\cdot\bm{q}\cos 2\phi+q^{2}\sin 2\phi,$ (29)
and $\mathcal{O}_{12}=\mathcal{O}_{21}$. The equation of motion, in Fourier space, can then be written as $\partial_{t}\delta\varphi_{a}=-\Gamma K_{ab}\delta\varphi_{b}$, where
$K_{ab}=q^{2}[(r+q^{2})\delta_{ab}+\alpha\epsilon_{ab}]+uA^{2}\mathcal{O}_{ab}.$ (30)
The stability of the traveling wave solution is given by the signs of the real parts of the eigenvalues of $K$. Modes with positive eigenvalues are stable, while modes with negative are unstable. For two-by-two matrices, the eigenvalues are given by the trace and determinant. We have
$\mathrm{Tr}\,K$ $=2q^{2}(r+q^{2})+x^{2}\mathrm{Tr}\,\mathcal{O},$ (31)
$\det K$ $=q^{4}\left[(r+q^{2})^{2}+\alpha^{2}\right]$ (32)
$\quad+q^{2}(r+q^{2})x^{2}\mathrm{Tr}\,\mathcal{O}+x^{4}\mathrm{det}\,\mathcal{O},$
where $x^{2}=uA^{2}$, and after some calculations, we obtain
$\mathrm{Tr}\,\mathcal{O}$ $=8k^{2}+4q^{2}$ (33)
$\det\mathcal{O}$ $=16[k^{4}+k^{2}q^{2}+(\bm{k}\cdot\bm{q})^{2}]+3q^{4}.$ (34)
Thus, we can calculate the eigenvalues of $K$, as follows
$\lambda_{\pm}$ $=q^{2}(r+q^{2})+2(2k^{2}+q^{2})x^{2}$ (35)
±$\quad\pm\sqrt{q^{4}(x^{4}-\alpha^{2})-(4x^{2}\bm{k}\cdot\bm{q})^{2}}.$
If we insert the amplitude criterion, $x^{2}=uA^{2}=|r|-k^{2}$ into the eigenvalue, we obtain
$\lambda_{\pm}$ $=q^{2}(|r|+q^{2}-2k^{2})+4k^{2}(|r|-k^{2})-2q^{2}k^{2}$ (36)
±$\quad\pm\sqrt{(|r|-k^{2})^{2}[q^{4}-(4\bm{k}\cdot\bm{q})^{2}]-\alpha^{2}q^{4}}.$
The most unstable mode is the one with a negative sign when the perturbation is orthogonal to the traveling wave. Furthermore, a non-zero value for $\alpha$ contributes to making the solution more stable. The most unstable mode is therefore obtained by setting $\alpha^{2}=\bm{k}\cdot\bm{q}=0$, which yields
$\lambda_{-}=q^{4}+4k^{2}(|r|-k^{2})-q^{2}k^{2}.$ (37)
@@
By assumption, $k^{2}<|r|$, so only the last term can make the solution unstable. If $k$ has the minimum value for the wavevector, $k_{0}=2\pi/L$ where $L$ is the linear system size, then $q^{2}\geq k^{2}$. Thus $q^{4}-q^{2}k^{2}\geq 0$ and $\lambda_{-}>0$. This shows that the long wavelength solution is always stable. The marginal case, when $|r|=k^{2}$, corresponds to the constant solution $\varphi_{a}=0$. There may exist solutions where $k=2\pi n/L$ for $n>1$, which corresponds to wrapping a line in the domain off $\varphi_{a}(\bm{x})$ several times around the center of the potential, but in the $\alpha=0$ case, at least, this is a metastable state.
@@ Appendix C Numerical methods
This has the disadvantage that high-frequency oscillations are not detected by $\nabla^{2}$ as it only measures sites at an even number of lattice sites away from the diagonal. We therefore apply a different operator, ~$\tilde{\nabla}^{2}$ with the stencil

### v3

@@ Appendix A Linear stability of $\bar{\varphi}=0$ solution
We now consider the stability of the traveling-wave solution of the $\mathrm{SO}(2)$ model. We write it in polar form as $\varphi_{a}(\bm{x},t)=A(\bm{k})[e^{\epsilon\phi(\bm{x},t;\bm{k})}]_{ab}\hat{\varphi}_{b}$, were $\hat{\varphi}$ is a unit vector. The solution in Eq. (18) corresponds to $\hat{\varphi}=(0,1)^{T}$, while a rotation of $\hat{\varphi}$ corresponds to a phase shift. The amplitude is given by $uA^{2}=-r-k^{2}$, and $\phi=\bm{k}\cdot\bm{x}-\alpha k^{2}t$. The wave-vector $\bm{k}$ thus points in the direction of travel. We then consider perturbations around this state, $A\rightarrow A+\rho$ and $\phi\rightarrow\phi+\theta$, to first order in the perturbations, We can find the stability of the solution by considering
$\varphi_{a}(\bm{x},t)$ $=[A(\bm{k})+\rho(\bm{x},t)][e^{\epsilon[\phi(\bm{x},t;\bm{k})+\theta(\bm{x},t)]}]_{ab}\hat{\varphi}_{a},$ (23)
$=[A(\bm{k})+\rho(\bm{x},t)][\cos\{\phi(\bm{x},t;\bm{k})+\theta(\bm{x},t)\}\delta_{ab}+\sin\{\phi(\bm{x},t;\bm{k})+\theta(\bm{x},t)\}\epsilon_{ab}]\hat{\varphi}_{a}.$ (24)
We insert this into the equation of motion Eq. (5), which gives
$[\partial_{t}\rho+\epsilon(-\alpha k^{2}\rho_{0}-\alpha k^{2}\rho+\rho_{0}\partial_{t}\theta)]e^{\epsilon(\phi+\theta)}\hat{\varphi}=\big\{[(r+u\rho_{0}^{2})+\alpha\epsilon][-\rho_{0}k^{2}+(a_{2}+\epsilon b_{2})\rho+(c_{2}+\epsilon d_{2})\theta],$ (25)
$\hskip 51.21495pt+2u\rho_{0}[\rho_{0}(\nabla^{2}-k^{2})\rho+\epsilon(2\rho_{0}\bm{k}\cdot\bm{\nabla})\rho]-\rho_{0}k^{4}+(a_{4}+\epsilon b_{4})\rho+(c_{4}+\epsilon d_{4})\theta\big\}e^{\epsilon(\phi+\theta)}\hat{\varphi}$, (26)
to leading order, where the derivative operators
$a_{2}$ $=\nabla^{2}-k^{2},$ $a_{4}$ $=-[(\nabla^{2}-k^{2})^{2}-4(\bm{k}\cdot\bm{\nabla})^{2}],$ (27)
$b_{2}$ $=2\bm{k}\cdot\bm{\nabla},$ $b_{4}$ $=-[4(\nabla^{2}-k^{2})\bm{k}\cdot\bm{\nabla}],$ (28)
$c_{2}$ $=-2\rho_{0}\bm{k}\cdot\bm{\nabla},$ $c_{4}$ $=\rho_{0}[4(\nabla^{2}-k^{2})\bm{k}\cdot\bm{\nabla}],$ (29)
$d_{2}$ $=\rho_{0}\nabla^{2},$ $d_{4}$ $=-\rho_{0}\left[(\nabla^{2}-2k^{2})\nabla^{2}-4(\bm{k}\cdot\bm{\nabla})^{2}\right],$ (30)
only affect $\rho$ and $\theta$ within the parenthesis, not the exponential functions. With this form, we can change to a rotating frame of reference where we write a set of two linear equations for the perturbations as follows
$\partial_{t}\rho$ $=[-k^{2}a_{2}-\alpha b_{2}-2(r+k^{2})(\nabla^{2}-k^{2})+a_{4}]\rho+[-k^{2}c_{2}-\alpha d_{2}+c_{4}]\theta,$ (31)
$\rho_{0}\partial_{t}\theta$ $=[\alpha a_{2}-k^{2}b_{2}+\alpha k^{2}-4u(r+k^{2})\bm{k}\cdot\bm{\nabla}+b_{4}]\rho+[\alpha c_{2}-k^{2}d_{2}+d_{4}]\theta.$ (32)
Note that all coefficients here are independent of space and time. The stability of the solution is then given by the eigenvalues of the corresponding matrix in Fourier space
$\lambda_{\pm}(\bm{q};\bm{k})=$ (33)
$\quad k^{4}-|r|\left(k^{2}+q^{2}\right)-4({\bm{k}\cdot\bm{q}})^{2}+2i\alpha({\bm{k}\cdot\bm{q}})-q^{4}\pm\big\{k^{8}+k^{4}\left(-4({\bm{k}\cdot\bm{q}})^{2}+q^{4}-4q^{2}|r|+|r|^{2}\right)$
$\quad+2k^{6}\left(q^{2}-|r|\right)+2k^{2}|r|\left(4({\bm{k}\cdot\bm{q}})^{2}-q^{4}+q^{2}|r|\right)+q^{4}(4({\bm{k}\cdot\bm{q}})-i\alpha)^{2}+4({\bm{k}\cdot\bm{q}})q^{2}|r|(4({\bm{k}\cdot\bm{q}})-i\alpha)+q^{4}|r|^{2}\big\}^{\frac{1}{2}}.$
Here, the condition $\mathrm{Re}(\lambda)>0$ implies that the corresponding eigen-mode grows exponentially. The most unstable eigenvalue is $\lambda_{+}$, which corresponds to perturbations parallel to the traveling wave ($\bm{k}\cdot\bm{q}=kq$). This mode is illustrated with solid lines in Fig. 8, for different values of $k$ and $\alpha$. For plots of orthogonal perturbations, $\bm{k}\cdot\bm{q}=0$, and the imaginary part, see the Supplemental Material [57]. We note that the real part has a zero when the wave-vector of the perturbation matches the traveling wave, i.e. $\lambda(\bm{q}=\bm{k};\bm{k})=i\alpha k^{2}$, which corresponds to a phase-shift. The $\nabla^{4}$ ensures the stability against high wave-number perturbations. We set $\bm{q}\cdot\bm{k}=qk$, and expand to obtain
$\lambda_{+,\parallel}(\bm{q};\bm{k})$ $=2i\alpha kq-D_{2,\parallel}(k)q^{2}+i\Omega_{\parallel}(k)q^{3}-D_{4,\parallel}(k)q^{4}+\mathcal{O}(q^{5}).$ (34)
@@
This approximate form is plotted as dashed lines in Fig. 8, and compared with solid lines that correspond to the full expression. The important parameter here is $D_{2,\parallel}=\frac{|r|-3k^{2}}{|r|-k^{2}}k^{2},$ which changes sign at the onset of the Eckhaus instability, thus requiring $k^{2}<\frac{1}{3}|r|$ as the condition for stability [29, 51]. This shows that the long wavelength solution is always stable. For a finite system of linear dimension $L$, there may exist solutions where $k=2\pi n/L$ for $n>1$, corresponding to wrapping a line in the domain off $\varphi_{a}(\bm{x})$ several times around the center of the potential, but in the $\alpha=0$ case, at least, this is a metastable state.
Figure 8: Eigenvalues for perturbation away from the traveling wave state, for different values of $k$, with $|r|=1$ and $\bm{k}\cdot\bm{q}=kq$, for different values of $k=|\bm{k}|$ and $\alpha$. The rightmost panel corresponds to a system beyond the Eckhaus-instability $k=1/\sqrt{3}\approx 0.577$.
@@ Appendix C Numerical methods
This has the disadvantage that high-frequency oscillations are not detected by $\nabla^{2}$ as it only measures sites at an even number of lattice sites away from the diagonal. We therefore apply a different operator, $\tilde{\nabla}^{2}$ with the stencil